OpenStudy (anonymous):

write 3logx-log2y+2logz as a single log

OpenStudy (anonymous):

\[\log{x^3}-\log{2y}+\log{z^2}=\log{\frac{x^3z^2}{2y}}\]

OpenStudy (anonymous):

\[3\log(x)=\log(x^3)\] \[\log(x^3)-log(2y)=\log(\frac{x^3}{2y})\] \[2\log(z)=\log(z^2)\] so you get \[\log(\frac{x^3z^2}{2y})\]

OpenStudy (anonymous):

I will compare this to the answer in the book real quick!

OpenStudy (anonymous):

Thank ya'll so much!! God bless u for your help!